OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, 9); Slope: 5

OpenStudy (anonymous):

A. y – 1 = 5(x + 9) B. y – 9 = 5(x + 1) C. y – 9 = 5(x – 1) D. y + 9 = 5(x – 1)

OpenStudy (anonymous):

whats with the question marks? kind of hard to tell what the problem is.

OpenStudy (anonymous):

i'm sorry it's my computer

OpenStudy (anonymous):

ok i see it now

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

if you need help with point slope this is a good site http://www.freemathhelp.com/point-slope.html

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ 9}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies 5 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (anonymous):

yup C

OpenStudy (anonymous):

thanks i have more

OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–3, 7); Slope: 4

OpenStudy (anonymous):

A. y – 7 = 4(x – 3) B. y – 7 = 4(x + 3) C. y + 7 = 4(x + 3) D. y + 7 = 4x – 12

OpenStudy (ilovepie):

The equation for point-slope form is \[y-y _{1}=m(x-x _{1})\] where m is the slope, y1 is the y-intercept, and x1 is the x-intercept. If the points are (-3,7) and the slope is 4, the equation in point-slope form is A. y-7=4(x-3)